OpenStudy (anonymous):

solve by completing the square x^2 -6x =27

OpenStudy (anonymous):

add 9 on both sides

OpenStudy (anonymous):

half of(-6\) is \(-3\) and \((-3)^2=9\) so you write \[(x-3)^2=27+9\] and go from there

OpenStudy (anonymous):

i meant " half of \(-6\) is \(-3\)"

OpenStudy (anonymous):

so 3,-9?????